# Liquidity fragmentation on decentralized exchanges

### v5

@@ Abstract
We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs). Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume. Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume. In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
@@
We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs). Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume. Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume. In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
@@
Liquidity fragmentation on decentralized exchanges
@@ 1 Introduction
Automated market makers such as Uniswap v3 provide a unique environment to investigate the market for liquidity. While there are various new institutional details that animate these exchanges, for our purposes two are economically important. First, in automated exchanges, liquidity demand and supply can easily be distinguished: users either supply or demand liquidity. Because of this, we can isolate the effect of transactions costs on each side of the market for liquidity. Second, prices are not set by market participants, but are automatically calculated as a function of liquidity demand and supply. Thus, costs are not passed from liquidity suppliers to demanders through prices and are easier to observe. These unique features allow us to investigate, theoretically and empirically, how transactions costs affect liquidity supply. Understanding how costs affect the market for liquidity is central to understanding how any asset prices are formed and whether asset exchange is efficient. Indeed, as Demsetz (1968) puts it “the question that is relevant for efficiency is whether or not the cost is appropriately economized.”
@@
Uniswap v3 provides two innovations over the previous v2. First, liquidity suppliers and demanders select into trading places (called pools) that differ on transaction fees. Each asset pair to be traded on up to four liquidity pools that only differ in the compensation for liquidity providers: in particular, liquidity fees can be equal to 1, 5, 30, or 100 basis points and the corresponding tick sizes are 1, 10, 60, or 200 basis points. These proportional fees are paid by liquidity demanders and are the only source of remuneration to liquidity suppliers. (These fees, as we discuss below, are similar to the make-take fees that are prevalent in limit order markets.) Second, on Uniswap v3, liquidity suppliers can submit “concentrated liquidity." Even though their liquidity is passively supplied, they can choose the price range over which it is supplied. With volatile assets, these concentrated liquidity positions can become stale and require rebalancing.
@@
We first present a simple model of liquidity supply in which there are two markets – one has high fees, and one has low fees. Liquidity demanders, who can either be small or large, choose a venue based on the fees each venue charges. The decisions of the liquidity demanders determine the payoff to the liquidity suppliers. Liquidity suppliers have heterogeneous endowments, interpretable as different capital constraints — low-endowment liquidity providers are akin to retail traders, whereas high-endowments stand in for large institutional investors or quantitative funds. Further, liquidity providers incur a fixed cost (i.e., gas price) each time they update their position. Traders demanding liquidity route their orders first to the low-fee pool to minimize transaction costs. The high-fee pool executes the residual order flow from large traders who exhaust all liquidity at the cheaper venue. As a result, low-fee markets are actively traded and require frequent liquidity updates whereas high-fee pools have a longer liquidity update cycle since they absorb fewer trades.
@@
We establish conditions under which there is fragmentation or consolidation. Specifically, even in this simple framework, there is a robust parameter range in which liquidity does not naturally concentrate on one of the exchanges. Both pools can attract a positive market share if the gas price is large enough due to economies of scale. Liquidity providers trade off a higher revenue per unit of time in the low-fee pool (driven by the larger trading volume) against the additional gas cost required for active liquidity management. As a result, liquidity provider clienteles emerge in equilibrium. Liquidity providers with large endowments gravitate towards low-fee markets, as they are best positioned to frequently update their position. In contrast, smaller market makers choose to passively provide liquidity on high-fee markets where they only trade against large orders being routed there. They optimally trade off a lower execution probability against higher fees per unit of volume and a lower liquidity management cost per unit of time.
@@
Not only does liquidity fragment, but it differs in both use and type across the two markets. A small number of highly active large liquidity providers, potentially institutional investors and hedge funds, primarily trade against numerous small incoming trades on pools with low fees. In contrast, high-fee pools involve less frequent trading between a substantial number of capital-constrained passive liquidity providers (e.g., retail market makers) on one side and a few sizeable incoming orders on the other. As the fixed gas fee affects liquidity providers pool choice, changes in the common fixed market access fee differentially affects the liquidity supply on the two pools. Specifically it reduces market quality (in the sense of lower posted liquidity) on the low fee pool.
@@
Our findings also indicate that liquidity fragmentation can enhance market quality, as measured by a lower implementation shortfall — a metric that combines both realized gains from trade as well as trading costs. In a single-pool market, selecting the optimal fee involves balancing these costs. A fee that is too low fails to attract liquidity providers with smaller endowments and thus more sensitive to fixed costs, leading to unrealized gains. Conversely, a very high fee results in prohibitively high trading costs. A two-pool market with heterogeneous fees provides enhanced flexibility, offering two instruments to independently manage the two components of the implementation shortfall. The higher fee determines the marginal liquidity provider LP entering the market, and therefore the realized gains from trade. The lower fee pool, by attracting LP with larger endowments, can reduce transaction costs. We demonstrate that a two-pool fee structure can always be designed to yield a lower implementation shortfall compared to any single-pool arrangement.
@@
We find robust evidence of liquidity supply clienteles across pools. The average liquidity deposit is 107.5% larger on the low-fee pool, after controlling for daily volume and return volatility. At the same time, high-fee pools’ market share is 21 percentage points higher. The results point to an asymmetric match between liquidity supply and demand: large liquidity suppliers are matched with small liquidity demanders on low-fee pools, whereas small liquidity suppliers trade with a few large orders on the high-fee pool.
We then turn to the common fixed cost of accessing the market, or gas fees. The degree of market fragmentation depends on the magnitude of gas costs on the Ethereum blockchain. In the model, a higher gas price leads to a shift in liquidity supply from the low- to the high-fee pool as active position management becomes relatively more costly for the marginal liquidity supplier. We find that a one standard deviation increase in gas prices corresponds to a 4.63 percentage points decrease in the low-fee pool market share, and a 29% drop in liquidity inflows on days when gas costs are elevated. As liquidity suppliers leave the low-fee pool, the time between position updates becomes 4.19% shorter since incoming order flow depletes liquidity at a faster pace. The effect reinforces the expected gas cost difference between the two pools, and more liquidity suppliers switch to the high-fee venue.
Finally, we examine differences in adverse selection costs across high- and low-fee pools. To do so, we use the loss-versus-rebalancing (LVR) price impact measure in Milinois, Moallemi, Roughgarden, and Zhang (2023), as well as impermanent loss (Heimbach, Schertenleib, and Wattenhofer, 2022). Our results reveal that low-fee pools face a greater permanent price impact, aligning with the notion that higher liquidity fees dissuade arbitrageurs. At the same time, arbitrage activity remains efficient: differences between high- and low-fee pool deviations from hourly centralized exchange prices are not significant. As a robustness check to our main analysis, we document that adverse selection is not a major driver of the relationship between gas prices and market fragmentation.
@@
We also relate to a rich literature on market fragmentation and differential fees. Closest to our paper, Battalio, Corwin, and Jennings (2016) and Cimon (2021) study the trade-off between order execution risk and compensation for liquidity provision in the context of make-take fee exchanges. However, Battalio, Corwin, and Jennings (2016) specifically addresses the issue of the broker-customer agency problem, whereas our study focuses on liquidity providers who trade on their own behalf. In traditional securities markets, make-take fees are contingent on trade execution and proportional to the size of the order. On the other hand, gas costs on decentralized exchanges are independent of order execution, highlighting the significance of economies of scale (lower proportional costs for larger liquidity provision orders) and dynamic liquidity cycles (managing the frequency of fixed cost payments). Strategic brokers in Cimon (2021) provide liquidity alongside exogenous market-makers in a static setting. We complement this approach by modelling network externalities inherent in the coordination problem of heterogeneous liquidity suppliers. In our dynamic setup, this allows us to pin down the equilibrium duration of liquidity cycles and the relative importance of gas fixed costs.
@@
Our paper relates to a nascent and fast-growing literature on the economics of decentralized exchanges. Many studies (e.g., Aoyagi, 2020; Aoyagi and Ito, 2021; Park, 2022) focus on the economics of constant-function automated market makers, which do not allow liquidity providers to set price limits. In this restrictive environment, Capponi and Jia (2021) argue that market makers have little incentives to update their position upon the arrival of news to avoid adverse selection, since pro-rata clearing gives an advantage to arbitrageurs. Lehar and Parlour (forthcoming) solve for the equilibrium pool size in a setting where liquidity providers fully internalize information costs without rushing to withdraw positions at risk of being sniped. We argue that on exchanges that allow for limit or range orders, the cost of actively managing positions becomes a first-order concern, as liquidity suppliers need to re-set the price limits once posted liquidity no longer earns fees. Our empirical result on economies of scale echoes the argument in Barbon and Ranaldo (2021), who compare transaction costs on centralized and decentralized exchanges and find that high gas prices imply that the latter only become competitive for transactions over US\$100,000. Hasbrouck, Rivera, and Saleh (2022) argue that liquidity suppliers require remuneration. We complement the argument by stating that high fees might be necessary for some liquidity providers to cover the fixed costs of managing their position. In line with our theoretical predictions, Caparros, Chaudhary, and Klein (2023) find that liquidity providers reposition their quotes more often on Uniswap V3 pools built on Polygon, which features substantially lower gas fees. Finally, Heimbach, Schertenleib, and Wattenhofer (2022) document that after accounting for price impact, concentrated liquidity on Uniswap v3 pools results in increased returns for sophisticated participants but losses for retail traders.
@@ 2 Model
Consider the following continuous time model of trade in a single token, T. The expected value of the token, $v>0$, is common knowledge. Two risk neutral trader types consummate trade in this market: a continuum of liquidity providers (LPs) and a continuum of liquidity takers (LTs). Trade occurs because market participants have heterogeneous private values for the asset. In particular, liquidity providers have no private value for the token, while liquidity takers value the token at $v\left(1+{\cal I}\Delta\right)$. Here $\Delta>0$ are the gains from trade and ${\cal I}$ is an indicator that takes on the value of $1$ if the taker buys and $-1$ if the taker sells. In what follows for expositional simplicity, as in Foucault, Kadan, and Kandel (2013), we focus on a one sided market in which liquidity takers act as buyers.
Liquidity providers differ in their endowments of the token. Each provider $i$ can supply at most $q_{i}$ of the token, where $q_{i}$ follows a truncated Pareto distribution on $\left[1,Q\right]$. So,
$\varphi\left(q\right)=\Big{(}\frac{Q}{Q-1}\Big{)}\frac{1}{q^{2}}\;\;\text{ for}q\in\left[1,Q\right].$ (1)
The right skew of the Pareto distribution captures the idea that there are many low-endowment liquidity providers such as retail traders, but few high-capital LPs such as sophisticated quantitative funds. Heterogeneity in LP size is captured by $Q$, where a larger $Q$ naturally corresponds to a larger dispersion of endowments. Given the endowment distribution, collectively LPs supply at most
$S=\int_{1}^{Q}q\varphi\left(q\right)\mathop{}\!\mathrm{d}q=\frac{Q}{Q-1}\log Q$ (2)
@@
There are two types of liquidity takers: small and large. Small LTs arrive at the market at constant rate $\theta\mathop{}\!\mathrm{d}t$, and each demand one unit of the token. The large LT arrival time follows a Poisson process with rate $\lambda>0$. Conditional on arrival, a large LT demands $\Theta$ units of the token, where $\Theta>S$. That is, the large LT liquidity demand exceeds the maximum liquidity supply. Let $D_{t}$ denote aggregate liquidity demand. Then,
$\mathop{}\!\mathrm{d}D_{t}=\theta\mathop{}\!\mathrm{d}t+\Theta\mathop{}\!\mathrm{d}J_{t}\left(\lambda\right),$ (3)
where $J_{t}\left(\lambda\right)$ is the Poisson arrival process.
Liquidity demanders and suppliers can interact in two liquidity pools in which token trade occurs against a numéraire asset (cash). We assume that the terms of trade are fixed, so that all trades occur at the expected value of the token, $v$. We do this by assuming prices in both pools satisfy a linear bonding curve, so there is no price impact of trade. [^4] In practice, most decentralized exchanges use convex bonding curves, for example constant product pricing. So, for a pool with $T$ tokens and $N$ of the numéraire good,
$vT+N=\text{constant}.$ (4)
It is well known that price impact costs can lead to order splitting and fragmentation. Fixing the terms of trade allows us to focus on how fees affect liquidity fragmentation.
Fees are levied on liquidity takers as a fraction of the value of the trade and distributed pro rata to liquidity providers. The pools have different fees. One pool charges a low fee, and one pool charges a high fee which we denote $\ell$ and $h$ respectively. Specifically, to purchase $\tau$ units of the token on the low fee pool, the total cost to a taker is $\tau\left(v+\ell\right)$. The LPs in the pool receive $\tau\ell$ in fees.
In addition, consistent with gas costs on Ethereum, any interaction with a liquidity pool (for example, trading or managing liquidity) incurs a fixed execution cost $\Gamma>0$. It is important to note that gas fees on decentralized exchanges differ from trading fees on traditional exchanges. First, they are not set by individual exchanges to compete with each other, but are a common transaction cost across trading venues. Second, they are levied on a per-order rather than per-share basis: this implies economies of scale for larger orders. Third, there is significant time variation in gas fees which will allow us to identify the impact of transaction costs on liquidity pool market shares. [^5] Empirical properties of gas fees are exhibited in Lehar and Parlour (forthcoming), while Capponi and Jia (2021), consider traders who affect the gas price.
To ensure that both small and large liquidity traders participate in the market, we assume that the gains from trade are larger than the aggregate transaction cost, including the pool fee and the gas price. To rule out trivial cases, we further assume that $Q\ell-\Gamma>0$. The condition ensures that there are at least some liquidity providers who can earn positive expected profit on the low fee pool.
@@
The gains from trade are sufficiently large so that all liquidity takers participate in the market, and the fixed costs are sufficiently low so that both pools can attract liquidity.
  1. i.
    Gains from trade are sufficiently large $\Delta>h+\frac{\Gamma}{v}$.
  2. ii.
    The low fee pool can attract liquidity $Q\ell-\Gamma>0$
We follow Foucault, Kadan, and Kandel (2013) and partition the continuous timeline into liquidity cycles. A liquidity cycle starts with an empty pool (zero liquidity offered) which triggers LP token deposits and ends when incoming trades deplete the liquidity supply. The first liquidity cycle starts at $t=0$, and the sequence of events within a cycle is as follows: each liquidity provider deposits their token into the high fee pool or the low fee pool and pays the gas cost $\Gamma$, or withdraw from the market. The LPs do not interact with the pool again until the their liquidity is consumed at some random time $\tilde{\tau}$ and the next cycle begins. Figure 1 illustrates the model timing.
Figure 1: Model timing

### v7

# Fragmentation and optimal liquidity supply on decentralized exchanges

@@ Abstract
We investigate how liquidity providers (LPs) choose between high- and low-fee trading venues, in the face of a fixed common gas cost. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply yet execute only 21% of volume. Large LPs dominate low-fee pools, frequently adjusting out-of-range positions in response to informed order flow. In contrast, small LPs converge to high-fee pools, accepting lower execution probabilities to mitigate adverse selection and liquidity management costs. Fragmented liquidity dominates a single-fee market, as it encourages more liquidity providers to enter the market, while fostering LP competition on the low-fee pool.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, adverse selection
@@
We investigate how liquidity providers (LPs) choose between high- and low-fee trading venues, in the face of a fixed common gas cost. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply yet execute only 21% of volume. Large LPs dominate low-fee pools, frequently adjusting out-of-range positions in response to informed order flow. In contrast, small LPs converge to high-fee pools, accepting lower execution probabilities to mitigate adverse selection and liquidity management costs. Fragmented liquidity dominates a single-fee market, as it encourages more liquidity providers to enter the market, while fostering LP competition on the low-fee pool.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, adverse selection
@@
Fragmentation and optimal liquidity supply on decentralized exchanges
@@ 1 Introduction
In addition to aggregating information, asset markets allow agents to exhaust private gains from trade. While there is a well developed literature on the informativeness of prices, less is known about if and when markets effectively exhaust all gains from trade. [^1] Gains from trade comprise an idiosyncratic private value for the underlying asset, but also idiosyncratic preferences for trade speed or “liquidity.”Agents’ idiosyncratic value for the underlying asset are plausibly determined by their portfolio positions, and therefore independent of the trading place. By contrast, their idiosyncratic preference for liquidity determines market quality. In this paper, we exploit the unique design of a decentralized exchange to shed light on the market for liquidity and show, theoretically and empirically, that market fragmentation can improve trading efficiency.
Automated market makers such as Uniswap v3 provide a unique environment to investigate the market for liquidity. While there are various new institutional details that animate these exchanges, for our purposes three are economically important. First, in automated exchanges, liquidity demand and supply can easily be distinguished: users either supply or demand liquidity. Because of this, we can isolate the effect of transactions costs on each side of the market for liquidity. Second, costs and benefits incurred by liquidity suppliers are easier to observe because prices are not set by market participants, but are automatically calculated as a function of liquidity demand and supply. Thus, liquidity suppliers are not compensated through price impact. Third, market participants are pseudo-anonymous so we can identify and document liquidity suppliers at a high frequency. These unique features allow us to investigate, theoretically and empirically, how transactions costs affect liquidity supply.
@@
Uniswap v3 provides two innovations over the previous v2. First, liquidity suppliers and demanders select into trading places (called pools) that differ on transaction fees. Each asset pair to be traded on up to four liquidity pools that only differ in the compensation for liquidity providers: in particular, liquidity fees can be equal to 1, 5, 30, or 100 basis points and the corresponding tick sizes are 1, 10, 60, or 200 basis points. These proportional fees are paid by liquidity demanders and are the only source of remuneration to liquidity providers. (These fees, as we discuss below, are similar to the make-take fees that are prevalent in limit order markets.) Second, on Uniswap v3, liquidity providers can submit “concentrated liquidity." Even though their liquidity is passively supplied, they can choose the price range over which it is supplied. With volatile assets, these concentrated liquidity positions can become stale and require rebalancing.
@@
We present a simple model with trade between liquidity suppliers and two types of liquidity demand. Consistent with the design of v3, liquidity suppliers chose a market and then place their liquidity into a band around the current value of the asset. The posted liquidity is subject to a bonding curve and hence generates a price impact cost for the liquidity demanders (we emphasize that this does not remunerate the liquidity suppliers). Liquidity suppliers have heterogeneous endowments, interpretable as different capital constraints — low-endowment liquidity providers are akin to retail traders, whereas high-endowments stand in for large institutional investors or quantitative funds. Trade occurs against these positions either because a liquidity demander arrives who has experienced a liquidity shock or because the value of the asset has changed and an arbitrageur adversely selects the passive liquidity supply. Collectively, the decisions of the liquidity demanders determine the payoff to the liquidity suppliers. After large private or common value trades, liquidity providers rebalance their positions; to do so, liquidity providers incur a fixed cost (i.e., gas price) each time they update their position.
Traders demanding liquidity face two types of costs: first, the fee associated with their chosen pool (low or high) and second, the price impact costs generated by the pool’s bonding curve and supplied liquidity. We find that traders route small orders exclusively to the low-fee pool to obtain the all-in lowest cost. In contrast, large traders split their orders across both low- and high-fee liquidity pools. As a result, low-fee markets are actively traded and require frequent liquidity updates whereas high-fee pools have a longer liquidity update cycle since they absorb fewer trades.
@@
We establish conditions under which there is fragmentation or consolidation. Specifically, even in this simple framework, there is a robust parameter range in which liquidity does not naturally concentrate on one of the exchanges. Both pools can attract a positive market share if liquidity providers face gas fees and the adverse selection costs are sufficiently low. Liquidity providers trade off a higher revenue per unit of time in the low-fee pool (driven by the larger trading volume) against higher adverse selection as well as the additional gas cost required for active liquidity management. As a result, liquidity provider clienteles emerge in equilibrium. Liquidity providers with large endowments gravitate towards low-fee markets, as they are best positioned to frequently update their position. In contrast, smaller market makers choose to passively provide liquidity on high-fee markets where they only trade against large orders being routed there. They optimally trade off a lower execution probability against higher fees per unit of volume, reduced adverse selection, as well as a lower liquidity management cost per unit of time.
@@
Not only does liquidity fragment, but it differs in both use and type across the two markets. A small number of highly active large liquidity providers, potentially institutional investors and hedge funds, primarily trade against numerous small incoming trades on pools with low fees. In contrast, high-fee pools involve less frequent trading between a substantial number of capital-constrained passive liquidity providers (e.g., retail market makers) on one side and a few sizeable incoming orders on the other. As the fixed gas fee affects liquidity providers pool choice, changes in the common fixed market access fee differentially affects the liquidity supply on the two pools. Specifically, it reduces market quality (in the sense of lower posted liquidity) on the low fee pool.
As we distinguish between liquidity demanders who are trading to exploit gains from trade and liquidity demanders who are arbitraging common value changes, we can decompose returns to liquidity providers, and show that adverse selection is higher on the low fee pool. Given that the low fee pool is populated with larger liquidity suppliers, this suggests that institutional traders bear price risk.
@@
Our findings indicate that liquidity fragmentation can enhance market quality, as measured by total gains from trade. In a single fee market, a fee that is too low fails to attract liquidity providers with smaller endowments and thus more sensitive to fixed costs, leading to unrealized gains. Conversely, a very high fee results in prohibitively high trading costs and deters trade. A two-pool market with heterogeneous fees offers two instruments to independently manage costs. The higher fee determines the marginal liquidity provider LP entering the market, and therefore the realized gains from trade. The lower fee pool, by attracting LP with larger endowments, can reduce transaction costs. We demonstrate that a two-pool fee structure can always be designed to yield higher gains from trade than any single-pool arrangement.
@@
We find robust evidence of liquidity supply clienteles across pools. The average liquidity deposit is 107.5% larger on the low-fee pool, after controlling for daily volume and return volatility. At the same time, high-fee pools’ market share is 21 percentage points higher. The results point to an asymmetric match between liquidity supply and demand: large liquidity providers are matched with small liquidity demanders on low-fee pools, whereas small liquidity providers trade with a few large orders on the high-fee pool.
We then turn to the common fixed cost of accessing the market, or gas fees. The market shares of the liquidity pools depend on the magnitude of gas costs on the Ethereum blockchain. In the model, a higher gas price leads to a shift in liquidity supply from the low- to the high-fee pool as active position management becomes relatively more costly for the marginal liquidity provider. We find that a one standard deviation increase in gas prices corresponds to a 4.63 percentage points decrease in the low-fee pool market share, and a 29% drop in liquidity inflows on days when gas costs are elevated.
Consistent with our model, we find that liquidity providers in low-fee pools earn higher fee yields but face increased adverse selection costs. Specifically, the daily fee yield is 2.03 basis points larger on low-fee pools. On the other hand, the permanent price impact as measured by loss-versus-rebalancing (LVR, as in Milinois, Moallemi, Roughgarden, and Zhang, 2023) is 6.39 basis points or 81% greater in low-fee pools compared to high-fee ones. However, despite this difference, the deviations in prices between high- and low-fee pools and those on centralized exchanges do not differ significantly.
@@
We also relate to a rich literature on market fragmentation and differential fees. Closest to our paper, Battalio, Corwin, and Jennings (2016) and Cimon (2021) study the trade-off between order execution risk and compensation for liquidity provision in the context of make-take fee exchanges. However, Battalio, Corwin, and Jennings (2016) specifically addresses the issue of the broker-customer agency problem, whereas our study focuses on liquidity providers who trade on their own behalf. In traditional securities markets, make-take fees are contingent on trade execution and proportional to the size of the order. On the other hand, gas costs on decentralized exchanges are independent of order execution, highlighting the significance of economies of scale (lower proportional costs for larger liquidity provision orders) and dynamic liquidity cycles (managing the frequency of fixed cost payments). Strategic brokers in Cimon (2021) provide liquidity alongside exogenous market-makers in a static setting. We complement this approach by modelling network externalities inherent in the coordination problem of heterogeneous liquidity providers. In our dynamic setup, this allows us to pin down the equilibrium duration of liquidity cycles and the relative importance of gas fixed costs.
@@
Our paper relates to a nascent and fast-growing literature on the economics of decentralized exchanges. Many studies (e.g., Aoyagi, 2020; Aoyagi and Ito, 2021; Park, 2022) focus on the economics of constant-function automated market makers, which do not allow liquidity providers to set price limits. In this restrictive environment, Capponi and Jia (2021) argue that market makers have little incentives to update their position upon the arrival of news to avoid adverse selection, since pro-rata clearing gives an advantage to arbitrageurs. Lehar and Parlour (forthcoming) solve for the equilibrium pool size in a setting where liquidity providers fully internalize information costs without rushing to withdraw positions at risk of being sniped. We argue that on exchanges that allow for limit or range orders, the cost of actively managing positions becomes a first-order concern, as liquidity providers need to re-set the price limits once posted liquidity no longer earns fees. Our empirical result on economies of scale echoes the argument in Barbon and Ranaldo (2021), who compare transaction costs on centralized and decentralized exchanges and find that high gas prices imply that the latter only become competitive for transactions over US\$100,000. Hasbrouck, Rivera, and Saleh (2022) argue that liquidity providers require remuneration. We complement the argument by stating that high fees might be necessary for some liquidity providers to cover the fixed costs of managing their position. In line with our theoretical predictions, Caparros, Chaudhary, and Klein (2023) find that liquidity providers reposition their quotes more often on Uniswap V3 pools built on Polygon, which features substantially lower gas fees. Finally, Heimbach, Schertenleib, and Wattenhofer (2022) document that after accounting for price impact, concentrated liquidity on Uniswap v3 pools results in increased returns for sophisticated participants but losses for retail traders.
@@ 2 Model
Asset and agents.
Consider a continuous time model of trade in a single token T with expected value $v_{t}>0$. Three risk neutral trader types consummate trade in this market: a continuum of liquidity providers (LPs), liquidity traders (LTs), and arbitrageurs (A). Trade occurs either because public news triggers a change in the common value of the asset, or because market participants have heterogeneous private values for the asset.
Arrival times of news and private value shocks follow independent Poisson processes with rates $\eta\in\left(0,1\right)$ and $1-\eta$, respectively. [^5] This is without loss of generality, as what matters in the model is the relative arrival rate of news relative to liquidity traders. For notational compactness, we first characterize the generic shock distribution and then describe its effects on arbitrageurs or liquidity traders. Conditional on an event at time $t$, the asset value changes to $v_{t}\left(1+{\cal I}\tilde{\delta}\right)$ for all traders in the case of a common value shock, or for an arriving (LT) in the case of the private value shock. Here, ${\cal I}$ is an indicator that takes on the value of $1$ if the taker buys and $-1$ if the taker sells. The value innovation $\tilde{\delta}$ has a probability density
$\phi\left(\delta\right)=\frac{1}{2\Delta\sqrt{1+\delta}}\;\;\text{ for }\delta\in\left[0,\Delta^{2}-1\right],$ (1)
thus $\sqrt{1+\tilde{\delta}}$ is uniformly distributed between $\left[1,\Delta\right]$. This assumption is innocuous and made for tractability purposes.
When news arrives, the innovation is to the common value of the token. (As we are agnostic as to the source of value of cryptocurrencies, this common value shock could include the possibility of resale on another exchange.) After such a shock, an arbitrageur A trades with the liquidity providers whenever profitable, and LPs face an adverse selection loss. Conversely, when a liquidity trader enters the market, they experience a private value shock — and liquidity providers continue to value the token at $v_{t}$. In what follows for expositional simplicity, as in Foucault, Kadan, and Kandel (2013), we focus on a one sided market in which liquidity takers act as buyers, and news lead to an increase in token value.
Liquidity providers (LP) differ in their endowments of the token. Each provider $i$ can supply at most $q_{i}\mathop{}\!\mathrm{d}i$ of the token, where $q_{i}$ follows an exponential distribution with scale parameter $\lambda$. The right skew of the distribution captures the idea that there are many low-endowment liquidity providers such as retail traders, but few high-capital LPs such as sophisticated quantitative funds. Heterogeneity in LP size is captured by $\lambda$, where a larger $\lambda$ naturally corresponds to a larger dispersion of endowments and larger aggregate liquidity supply. Given the endowment distribution, collectively LPs supply at most
$S=\int_{0}^{\infty}q_{i}\frac{1}{\lambda}e^{-\frac{q_{i}}{\lambda}}\mathop{}\!\mathrm{d}i=\lambda$ (2)
@@
Trading environment.
Traders can interact in two liquidity pools in which token trade occurs against a numéraire asset (cash). At the start of the trading game, each liquidity provider (LP) deposits liquidity to a single pool within a symmetric price band around the current asset value $\left[\frac{v}{\left(1+r\right)^{2}},v(1+r)^{2}\right]$, where $r\geq 0$. Here, we make use of the fact that V3 features “price bands,” and thus liquidity can be consumed with a bounded price impact. Within this range, prices in both pools satisfy a constant product bonding curve as in Adams, Zinsmeister, Salem, Keefer, and Robinson (2021). In particular, for pool $k$,
$\underbrace{\left(T_{k}+\frac{L_{k}}{\sqrt{v}\left(1+r\right)}\right)}_{\text{virtual token reserves}}\underbrace{\left(T_{k}v+L_{k}\frac{\sqrt{v}}{1+r}\right)}_{\text{virtual numeraire reserves}}=L_{k}^{2},$ (3)
where $T_{k}$ is the amount of tokens deposited on pool $k$ and $L_{k}$ is the liquidity level of pool $k$, defined as
$L_{k}=\frac{T_{k}}{\frac{1}{\sqrt{v}}-\frac{1}{\sqrt{v}}\left(1+r\right)}.$ (4)
To purchase $\tau$ tokens, a trader needs to deposit an amount $n\left(\tau\right)=\tau T_{k}\frac{v\left(1+r\right)}{\tau+\left(1+r\right)\left(T_{k}-\tau\right)}$ of numéraire into the pool, where $n\left(\tau\right)$ is the solution to the invariance condition
$\underbrace{\left(T_{k}-\tau+\frac{L_{k}}{\sqrt{v}\left(1+r\right)}\right)}_{\text{virtual token reserves}}\underbrace{\left(T_{k}v+n\left(\tau\right)+L_{k}\frac{\sqrt{v}}{1+r}\right)}_{\text{virtual numeraire reserves}}=L_{k}^{2}.$ (5)
Fees are levied on liquidity takers as a fraction of the value of the trade and distributed pro rata to liquidity providers. Crucially, the pools have different fees. One pool charges a low fee, and one pool charges a high fee which we denote $\ell$ and $h$ respectively. Specifically, to purchase $\tau$ units of the token on the low fee pool, the total cost to a taker is $\left(1+\ell\right)n\left(\tau,T_{\ell}\right)$. The LPs in the pool receive $\ell n\left(\tau,T_{\ell}\right)$ in fees. In addition, consistent with gas costs on Ethereum, all traders incur a fixed execution cost $\Gamma\mathop{}\!\mathrm{d}i>0$ to interact with the market.
Figure 1 illustrates the timing of the model.
Pool HPool LLPs deposit $q_{i}$in pool $k\in\left\{L,H\right\}$Small LTSmall LTLarge LTNews $\tilde{\delta}\in\left(\ell,h\right)$A trades on pool $L$LP rebalance on pool $L$Large LTSmall LTNews $\tilde{\delta}>h$A trades on both poolsLP rebalance on both pools
Figure 1: Model timing
To ensure the possibility of liquidity re-balancing in both pools, we assume that innovations are large enough to ensure that LPs may need to rebalance their position on the high fee pool or:
@@
The size of innovations are sufficiently large so that there is a positive probability that liquidity providers need to re-balance on the high-fee pool. That is, $\Delta>\left(1+r\right)\sqrt{1+h}$.